Excel to do the calculations. Be sure to write clear steps for your derivations.5. (25 points) After extensive medical and marketing research, Pill, Inc., believes it can penetrate the pain reliever market
Answer #1

Headache only Pill Selling Price after Discount = $8.25 - (7/100 * $8.25) = 7.6725  
Total Units to be sold of Headache only Pill = 3 Million
Headache only Pill Cost for the 1st year = $4.15
Headache only Pill Cost for the 2nd Year = $4.15 * 1.03 (Inflation) = $4.27
Headache only Pill Cost for the 3rd Year = $4.15 * 1.03 * 1.03 = $4.40
Income for the 1st Year = ($7.6725 * 3 Million) - ($4.15 * 3 Million) = $10.57 Million
Income for the 2nd Year = ($7.6725 * 3 Million) - ($4.27 * 3 Million) = $10.20 Million
Income for the 3rd Year = ($7.6725 * 3 Million) - ($4.40 * 3 Million) = $9.82 Million
Total Income = ($10.57 Million + $10.20 Million + $9.82 Million ) - Cost of Machinery($22 Million) = $8.59 Million
Tax on Income of 34% = ($8.59) * (34/100 * $8.59 Million)
Net Income = $5.67 Million

Headache and arthritis Pill Selling Price = $8.25 - (7/100 * $8.25) = 7.6725
Total Units to be sold of Headache and arthritis Pill = 4.5 Million
Headache and arthritis Pill Cost for the 1st year = $4.55
Headache and arthritis Pill Cost for the 2nd Year = $4.55*1.03 = $4.69
Headache and arthritis Pill Cost for the 3rd Year = $4.55*1.03*1.03 = $4.83
Income for the 1st Year = ($7.6725 * 4.5 Million) - ($4.55 * 4.5 Million) = $14.05 Million
Income for the 2nd Year = ($7.6725 * 4.5 Million) = ($4.69 * 4.5 Million) = $13.43 Million
Income for the 3rd Year = ($7.6725 * 4.5 Million) = (4.83 * 4.5 Million) = $12.79 Million
Total Income = ($14.05 Million + $13.43 Million + $12.79 Million) - Cost of Machinery ($32 Million) + Resale Value of Machine ($1 Million) = $9.27 Million
Tax on Income of 34% = $9.27 Million * (34/100 * 9.27 Million)
Net Income = $6.12 Million

So the company should produce Headache and arthritis Pill as the profit is Higher as Shown in the above Calculations.